# UStyle: Waterbody Style Transfer of Underwater Scenes by Depth-Guided Feature Synthesis

### v2

@@ V-A Baseline Models and Metrics
We use six SOTA models from artistic and photorealistic STx domains as baselines for performance evaluation. Our comparison includes artistic STx methods StyTR-2 [115], ASTMAN [17], MCCNet [16], and IEContraAST [116], alongside the photorealistic method PhotoWCT2 [4]. Additionally, we include AquaFuse [1], a physics-based underwater waterbody style transfer method, to benchmark our approach further. For a fair comparison, all models are fine-tuned on the UF7D dataset; specifically, their pre-trained weights are fine-tuned over $160$ K iterations with a batch size of $8$ while keeping all other hyperparameters at their recommended settings. For evaluation, we define four cross-domain underwater style transfer test cases: Blue (B) $\rightarrow$ Green (G), Green (G) $\rightarrow$ Blue (B), Deep Blue (DB) $\rightarrow$ Deep Green (DG), and Deep Green (DG) $\rightarrow$ Deep Blue (DB). Each case consists of $25$ content images and $25$ style images, generating $25\times 25=625$ unique test scenes per case, resulting in a total of $4\times 625=2,500$ test scenes.
Evaluation Metrics. We use five standard metrics to evaluate the performance of UStyle and other baseline methods. These complementary metrics assess both the structural preservation and perceptual quality of the stylized images generated by each model. The metrics are:
  • Root Mean Square Error (RMSE) [117] quantifies pixel-level accuracy between stylized and content images as: ${\tt{RMSE}}=\sqrt{\tt{MSE}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(x_{i}-y_{i})^{2}}$; where $x_{i}$ and $y_{i}$ are corresponding pixel values and $N$ is total number of pixels. The squared difference term amplifies larger errors, making RMSE especially sensitive to outliers and structural distortions.
  • Peak Signal-to-Noise Ratio (PSNR) [22] measures reconstruction quality as ${\tt{PSNR}}=10\cdot\log_{10}\left(\frac{L^{2}}{\tt{MSE}}\right)$; where $L=255$ represents the maximum possible pixel value in 8-bit images, and MSE is the mean squared error. PSNR’s logarithmic scale and dependency on $L$ make it particularly sensitive to illumination changes and global color shifts.
  • Structural Similarity Index (SSIM) [118] evaluates the geometric consistency of the generated image based on luminance, contrast, and changes in structural information. It is calculated as:
    ${\tt{SSIM}}(x,y)=\frac{(2\mu_{x}\mu_{y}+C_{1})(2\sigma_{xy}+C_{2})}{(\mu_{x}^{2}+\mu_{y}^{2}+C_{1})(\sigma_{x}^{2}+\sigma_{y}^{2}+C_{2})}$ (10)
    where $\mu$ and $\sigma$ represent mean and variance respectively, and $C_{1}$, $C_{2}$ are stability constants. The multiplicative combination of luminance ($\mu$), contrast ($\sigma$), and structure ($\sigma_{xy}$) terms makes SSIM particularly effective for evaluating perceptual image quality.
  • Gradient Magnitude Similarity Deviation (GMSD) [119] assesses edge preservation by exploiting image gradient information as: ${\tt{GMSD}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(g_{i}-\mu_{g})^{2}}$; where $g_{i}$ represents gradient magnitude similarity at pixel $i$, and $\mu_{g}$ is the mean. GMSD is particularly effective at detecting structural and textural variations while being relatively insensitive to global illumination changes.
  • Learned Perceptual Image Patch Similarity (LPIPS) [120] measures perceptual similarity as:
    ${\tt{LPIPS}}(x,y)=\sum_{l}\frac{1}{H_{l}\,W_{l}}\sum_{h,w}\Big{\|}w_{l}\odot(\phi_{l}(x)_{h,w}-\phi_{l}(y)_{h,w})\Big{\|}_{2}^{2}$ (11)
    Here, $\phi_{l}$ represents feature maps from layer $l$ of a pre-trained network, and $w_{l}$ are learned weights that calibrate the importance of different network layers. This metric captures high-level perceptual similarities that often correlate better with human perceptual judgment than traditional metrics.
This combination of metrics enables the evaluation of both low-level image statistics and high-level perceptual qualities. All evaluations are conducted on images normalized to $640\times 480$ resolution using their official metric implementations.
@@ V-B Qualitative Results
Figure 8: Qualitative comparison of underwater image style transfer (STx) results across different methods. The first two columns show content and style images, followed by stylized outputs from StyTR-2 [7], ASTMAN [17], MCCNet [6], IEContraAST [5], PhotoWCT2 [4], AquaFuse [1], and UStyle. Each row shows two examples of: B $\rightarrow$ G, G $\rightarrow$ B, DB $\rightarrow$ DG, and DG $\rightarrow$ DB transformations.
@@ V-C Quantitative Results
We conduct a comprehensive quantitative evaluation using five complementary metrics across various underwater STx scenarios from the UF7D test set. As presented in Table I(a), UStyle achieves 4×4$\times$4 × to 6×6$\times$6 × lower RMSE rates across all scenarios compared to SOTA artistic and photorealistic STx methods. It also consistently outperforms competing approaches in PSNR and SSIM, as shown in Table I(b). The closest competitor, AquaFuse, requires a carefully selected reference image and prior information such as incident angle and attenuation characteristics. In contrast, UStyle achieves improvements of $15\%$ (RMSE), $5.29\%$ (PSNR), and $5.71\%$ (SSIM) over AquaFuse while operating without any reference or prior knowledge — demonstrating its robustness and accuracy for waterbody STx.
@@
TABLE I: Quantitative comparison of averaged RMSE ($\downarrow$), PSNR ($\uparrow$), and SSIM ($\uparrow$) across different underwater STx methods are conducted on $2,500$ test samples in UF7D. The best score is highlighted in bold blue, while the second-best is shown in blue.
@@
ASTMAN [17] $52.29$ $54.53$ $48.91$ $63.74$
IEContraAST [116] $47.32$ $50.54$ $58.17$ $69.49$
MCCNet [16] $44.64$ $53.12$ $46.04$ $60.71$
PhotoWCT2 [4] $34.68$ $34.80$ $31.89$ $26.66$
StyTR-2 [115] $37.08$ $40.61$ $47.18$ $54.01$
AquaFuse [1] $10.45$ $10.19$ $10.15$ $10.24$
UStyle $9.39$ $9.35$ $9.50$ $8.96$
@@
ASTMAN [17] $14.95$ $14.37$ $14.99$ $13.16$
IEContraAST [116] $15.90$ $14.40$ $13.82$ $12.10$
MCCNet [16] $15.81$ $14.01$ $15.63$ $13.35$
PhotoWCT2 [4] $19.75$ $18.70$ $18.95$ $20.05$
StyTR-2 [115] $17.86$ $16.67$ $15.59$ $14.58$
AquaFuse [1] $27.77$ $28.00$ $28.03$ $27.94$
UStyle $28.73$ $28.72$ $28.64$ $29.14$
@@
ASTMAN [17] $0.8026$ $0.6334$ $0.8150$ $0.6397$
IEContraAST [116] $0.7641$ $0.6304$ $0.7887$ $0.6033$
MCCNet [16] $0.8095$ $0.6302$ $0.8290$ $0.6570$
PhotoWCT2 [4] $0.8675$ $0.7819$ $0.8324$ $0.8030$
StyTR-2 [115] $0.8494$ $0.6937$ $0.8199$ $0.7125$
AquaFuse [1] $0.9337$ $0.8708$ $0.7394$ $0.8812$
UStyle $0.9613$ $0.8818$ $0.8389$ $0.9126$
(b) (d) Average RMSE scores are shown on (top barchart) and the PSNR--SSIM scores are visualized in a 2D grid (bottom); evident from these results, UStyle achieves better overall performance than existing STx methods.
TABLE II: Quantitative comparison of averaged GMSD ($\downarrow$) and LPIPS ($\downarrow$) across different underwater STx methods, conducted on $2,500$ test samples in UF7D. The best score is highlighted in bold blue, while the second-best is shown in blue.
@@
ASTMAN [17] $0.0693$ $0.0318$ $0.0508$ $0.0179$
IEContraAST [116] $0.0694$ $0.0300$ $0.0556$ $0.0187$
MCCNet [16] $0.0262$ $0.0107$ $0.0386$ $0.0386$
PhotoWCT2 [4] $0.0622$ $0.0725$ $0.0762$ $0.0386$
StyTR-2 [115] $0.0551$ $0.0287$ $0.0380$ $0.0187$
AquaFuse [1] $0.0252$ $0.0631$ $0.0201$ $0.0471$
UStyle $0.0227$ $0.0564$ $0.0200$ $0.0431$
@@
ASTMAN [17] $0.2992$ $0.4849$ $0.3769$ $0.4091$
IEContraAST [116] $0.2762$ $0.4835$ $0.3896$ $0.4518$
MCCNet [16] $0.2838$ $0.4976$ $0.3677$ $0.3989$
PhotoWCT2 [4] $0.1676$ $0.2618$ $0.2409$ $0.1755$
StyTR-2 [115] $0.2372$ $0.3612$ $0.3624$ $0.3261$
AquaFuse [1] $0.4274$ $0.3902$ $0.5597$ $0.3789$
UStyle $0.1926$ $0.2235$ $0.1342$ $0.1373$
@@ V-D Ablation Studies
Additionally, we investigate the contribution of different loss components by evaluating three configurations of our model: (1) UStyle without clip loss (w/o CLIP), (2) UStyle without additional regularization losses, $L3=\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}$ (w/o $L3$), and (3) UStyle with all loss components. Table III presents the quantitative comparison across PSNR, SSIM, and RMSE for four underwater style transfer scenarios. The results show that the full UStyle model consistently performs better, demonstrating the effectiveness of depth supervision and loss regularization. From a quantitative perspective, depth supervision leads to a substantial improvement in all three metrics. The inclusion of additional losses further refines the model’s ability to balance structural preservation and perceptual realism. The full UStyle model achieves the highest PSNR and SSIM values while maintaining the lowest RMSE across all scenarios.
To further examine the impact of DA-WCT stylization, we employ Grad-CAM [121] visualizations to highlight the regions of interest in our model’s decision-making process. As illustrated in Fig. 10, depth supervision enables UStyle to focus more effectively on salient structures, ensuring spatially coherent style adaptation. The middle row displays activation maps generated with depth guidance, where attention is evenly distributed across foreground and background elements, enhancing structure retention. In contrast, the bottom row—corresponding to the model without depth supervision—reveals inconsistent attention patterns, leading to fine-detail loss and exaggerated distortions. Notably, DA-WCT ensures that complex structures, such as marine organisms and underwater artifacts, are better preserved by aligning style fusion with content depth information. These visualizations further validate the importance of incorporating depth guidance, reinforcing its role in improving geometric consistency and reducing artifacts in waterbody STx.
@@
TABLE III: For ablation study, we compare the performance of three variants: UStyle without clip loss (w/o CLIP), UStyle without additional regularization losses $L3$ (w/o $L3$), and UStyle with all loss components. Each of these variants is evaluated with WCT/DA-WCT on PSNR (↑), SSIM (↑), and RMSE (↓) metrics.
Data B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ G DG $\rightarrow$ B
Metric Model WCT DA-WCT WCT DA-WCT WCT DA-WCT WCT DA-WCT
PSNR 1: UStyle (w/o CLIP) $28.03$ $28.69$ $28.24$ $28.69$ $27.85$ $28.00$ $28.39$ $29.09$
2: UStyle (w/o $L3$) $27.96$ $28.36$ $28.21$ $28.64$ $27.84$ $27.91$ $28.43$ $28.97$
3: UStyle (Full) $28.09$ $28.73$ $28.24$ $28.72$ $27.91$ $28.64$ $28.43$ $29.14$
SSIM 1: UStyle (w/o CLIP) $0.9319$ $0.9549$ $0.8090$ $0.8758$ $0.6365$ $0.8046$ $0.8421$ $0.9102$
2: UStyle (w/o $L3$) $0.9337$ $0.9590$ $0.8004$ $0.8675$ $0.6087$ $0.7780$ $0.8388$ $0.9069$
3: UStyle (Full) $0.9349$ $0.9613$ $0.8260$ $0.8818$ $0.6727$ $0.8389$ $0.8506$ $0.9126$
RMSE 1: UStyle (w/o CLIP) $10.15$ $9.43$ $9.90$ $9.40$ $10.37$ $10.19$ $9.75$ $9.01$
2: UStyle (w/o $L3$) $10.23$ $9.77$ $9.93$ $9.45$ $10.38$ $10.28$ $9.70$ $9.15$
3: UStyle (Full) $10.08$ $9.39$ $9.89$ $9.35$ $10.31$ $9.50$ $9.70$ $8.96$

### v3

@@ V-A Baseline Models and Metrics
Seven state-of-the-art (SOTA) models from artistic and photorealistic STx domains are used for comparison, including learning-based models: StyTR-2 [deng2021stytr2], ASTMAN [deng2020arbitrary], MCCNet [wang2020arbitrary], IEContraAST [qi2020iecontrasast], PhotoWCT2 [chiu2022photowct2], StyleID [chung2024style], and the underwater physics-based method: AquaFuse [siddique2024aquafuse]. All models are fine-tuned on UF7D for $160$ K iterations with a batch size of $8$, using their default hyperparameters for fairness. For evaluation, we define four cross-domain underwater style transfer test cases: Blue (B) $\rightarrow$ Green (G), Green (G) $\rightarrow$ Blue (B), Deep Blue (DB) $\rightarrow$ Deep Green (DG), and Deep Green (DG) $\rightarrow$ Deep Blue (DB). Each case consists of $25$ content images and $25$ style images, yielding $25\times 25=625$ unique test scenes per case, for a total of $4\times 625=2,500$ test scenes.
Evaluation Metrics. We evaluate UStyle and all baseline methods using five standard image-quality metrics. These are: (i) RMSE [chai2014root] for pixel-level deviation; (ii) PSNR [hore2010image] for reconstruction fidelity; (iii) SSIM [wang2004image] to evaluate luminance, contrast, and structural consistency; (iv) GMSD [xue2014gradient] for edge preservation and local gradient similarity; and (v) LPIPS [zhang2018unreasonable] to capture high-level perceptual similarity using deep feature embeddings. Together, these metrics provide a comprehensive assessment of both structural accuracy and perceptual quality, with all results computed on images normalized to $640\times 480$ using official metric implementations.
Figure 6: Qualitative comparison of underwater image style transfer (STx) results across different methods. The first two columns show content and style images, followed by stylized outputs from StyTR-2 [deng2022stytr2], ASTMAN [deng2020arbitrary], MCCNet [deng2021arbitrary], IEContraAST [chen2021artistic], PhotoWCT2 [chiu2022photowct2], StyleID [chung2024style], AquaFuse [siddique2024aquafuse], and UStyle. Each row shows two examples of: B $\rightarrow$ G, G $\rightarrow$ B, DB $\rightarrow$ DG, and DG $\rightarrow$ DB transformations.
@@ V-C Quantitative Results
We conduct a comprehensive quantitative evaluation using five complementary metrics across various underwater STx scenarios from the UF7D test set. As presented in Table I(a), UStyle achieves $4\times$ to $6\times$ lower RMSE rates across all scenarios compared to SOTA artistic and photorealistic STx methods. It also consistently outperforms competing approaches in PSNR and SSIM, as shown in Table I(b). The closest competitor, AquaFuse, requires a carefully selected reference image and prior information such as incident angle and attenuation characteristics. In contrast, UStyle achieves improvements of $15\%$ (RMSE), $5.29\%$ (PSNR), and $5.71\%$ (SSIM) over AquaFuse while operating without any reference or prior knowledge — demonstrating its robustness and accuracy for waterbody STx.
@@
TABLE I: Quantitative comparison of averaged RMSE ($\downarrow$), PSNR ($\uparrow$), and SSIM ($\uparrow$) across different underwater STx methods are conducted on $2,500$ test samples in UF7D. The best score is highlighted in bold blue, while the second-best is shown in blue. See Table II for GMSD and LPIPs evaluation.
@@
ASTMAN [deng2020arbitrary] $52.29$ $54.53$ $48.91$ $63.74$
IEContraAST [qi2020iecontrasast] $47.32$ $50.54$ $58.17$ $69.49$
MCCNet [wang2020arbitrary] $44.64$ $53.12$ $46.04$ $60.71$
PhotoWCT2 [chiu2022photowct2] $34.68$ $34.80$ $31.89$ $26.66$
StyTR-2 [deng2021stytr2] $37.08$ $40.61$ $47.18$ $54.01$
StyleID [chung2024style] 48.44 53.90 55.05 52.55
AquaFuse [siddique2024aquafuse] $10.45$ $10.19$ $10.15$ $10.24$
UStyle (ours) 9.39 9.35 9.50 8.96
@@
ASTMAN [deng2020arbitrary] $14.95$ $14.37$ $14.99$ $13.16$
IEContraAST [qi2020iecontrasast] $15.90$ $14.40$ $13.82$ $12.10$
MCCNet [wang2020arbitrary] $15.81$ $14.01$ $15.63$ $13.35$
PhotoWCT2 [chiu2022photowct2] $19.75$ $18.70$ $18.95$ $20.05$
StyTR-2 [deng2021stytr2] $17.86$ $16.67$ $15.59$ $14.58$
StyleID [chung2024style] 14.72 13.73 14.32 14.62
AquaFuse [siddique2024aquafuse] $27.77$ $28.00$ $28.03$ $27.94$
UStyle (ours) 28.73 28.72 28.64 29.14
@@
ASTMAN [deng2020arbitrary] $0.8026$ $0.6334$ $0.8150$ $0.6397$
IEContraAST [qi2020iecontrasast] $0.7641$ $0.6304$ $0.7887$ $0.6033$
MCCNet [wang2020arbitrary] $0.8095$ $0.6302$ $0.8290$ $0.6570$
PhotoWCT2 [chiu2022photowct2] $0.8675$ $0.7819$ $0.8324$ $0.8030$
StyTR-2 [deng2021stytr2] $0.8494$ $0.6937$ $0.8199$ $0.7125$
StyleID [chung2024style] 0.3528 0.3503 0.2800 0.3527
AquaFuse [siddique2024aquafuse] $0.9337$ $0.8708$ $0.7394$ $0.8812$
UStyle (ours) 0.9613 0.8818 0.8389 0.9126
(a) (d) Average RMSE scores are shown on (top barchart) and the PSNR-SSIM scores are visualized in a 2D grid (bottom); evident from these results, UStyle achieves better overall performance than existing STx methods.
TABLE II: Quantitative comparison of GMSD ($\downarrow$) and LPIPS ($\downarrow$) across different underwater STx methods, averaged over $2,500$ test samples in UF7D.
@@
ASTMAN [deng2020arbitrary] $0.0693$ $0.0318$ $0.0508$ 0.0179
IEContraAST [qi2020iecontrasast] $0.0694$ $0.0300$ $0.0556$ $0.0187$
MCCNet [wang2020arbitrary] $0.0262$ 0.0107 $0.0386$ $0.0386$
PhotoWCT2 [chiu2022photowct2] $0.0622$ $0.0725$ $0.0762$ $0.0386$
StyTR-2 [deng2021stytr2] $0.0551$ $0.0287$ $0.0380$ $0.0187$
StyleID [chung2024style] 0.0710 0.0321 0.0685 0.0403
AquaFuse [siddique2024aquafuse] $0.0252$ $0.0631$ $0.0201$ $0.0471$
UStyle (ours) 0.0227 $0.0564$ 0.0200 $0.0431$
@@
ASTMAN [deng2020arbitrary] $0.2992$ $0.4849$ $0.3769$ $0.4091$
IEContraAST [qi2020iecontrasast] $0.2762$ $0.4835$ $0.3896$ $0.4518$
MCCNet [wang2020arbitrary] $0.2838$ $0.4976$ $0.3677$ $0.3989$
PhotoWCT2 [chiu2022photowct2] 0.1676 $0.2618$ $0.2409$ $0.1755$
StyTR-2 [deng2021stytr2] $0.2372$ $0.3612$ $0.3624$ $0.3261$
StyleID [chung2024style] 0.5164 0.4091 0.3710 0.4618
AquaFuse [siddique2024aquafuse] $0.4274$ $0.3902$ $0.5597$ $0.3789$
UStyle (ours) $0.1926$ 0.2235 0.1342 0.1373
@@
TABLE III: For ablation on each loss term’s contributions, we compare the performance of UStyle without: $\mathcal{L}_{ssim}$, $\mathcal{L}_{color}$, $\mathcal{L}_{fft}$, $\mathcal{L}_{percept}$, $\mathcal{L}_{clip}$ and ${L}_{3}$, compared to the full model. Each of these variants is evaluated with WCT/DA-WCT on PSNR (↑), SSIM (↑), and RMSE (↓) metrics.
Data B $\rightarrow$ G G $\rightarrow$ B DB $\rightarrow$ G DG $\rightarrow$ B
Metric Model WCT DA-WCT WCT DA-WCT WCT DA-WCT WCT DA-WCT
PSNR 1: UStyle (w/o $\mathcal{L}_{SSIM}$) $27.48$ $28.47$ $27.84$ $28.65$ $27.37$ $27.73$ $27.85$ $28.70$
2: UStyle (w/o $\mathcal{L}_{color}$ ) $27.58$ $28.62$ $27.96$ $28.74$ $27.73$ $28.22$ $27.92$ $28.76$
3: UStyle (w/o $\mathcal{L}_{fft}$ ) $27.43$ $28.56$ $27.56$ $28.15$ $27.29$ $27.62$ $28.02$ $28.84$
4: UStyle (w/o $\mathcal{L}_{percept}$) $27.93$ $28.76$ $28.29$ $28.80$ $27.49$ $27.93$ $28.15$ $28.95$
5: UStyle (w/o $\mathcal{L}_{clip}$) $28.03$ $28.69$ $28.24$ $28.69$ $27.85$ $28.00$ $28.39$ $29.09$
6: UStyle (w/o $L3$) $27.96$ $28.36$ $28.21$ $28.64$ $27.84$ $27.91$ $28.43$ $28.97$
7: UStyle (Full) $28.09$ $28.73$ $28.24$ $28.72$ $27.91$ $28.64$ $28.43$ $29.14$
SSIM 1: UStyle (w/o $\mathcal{L}_{SSIM}$) $0.9285$ $0.9434$ $0.7975$ $0.8693$ $0.5990$ $0.7771$ $0.8402$ $0.8981$
2: UStyle (w/o $\mathcal{L}_{color}$) $0.9307$ $0.9594$ $0.8001$ $0.8770$ $0.6157$ $0.7961$ $0.8370$ $0.9002$
3: UStyle (w/o $\mathcal{L}_{fft}$ ) $0.9291$ $0.9516$ $0.7990$ $0.8688$ $0.6014$ $0.7912$ $0.8391$ $0.9012$
4: UStyle (w/o $\mathcal{L}_{percept}$) $0.9309$ $0.9540$ $0.8102$ $0.8775$ $0.6216$ $0.8059$ $0.8456$ $0.9080$
5: UStyle (w/o $\mathcal{L}_{clip}$) $0.9319$ $0.9549$ $0.8090$ $0.8758$ $0.6365$ $0.8046$ $0.8421$ $0.9102$
6: UStyle (w/o $L3$) $0.9337$ $0.9590$ $0.8004$ $0.8675$ $0.6087$ $0.7780$ $0.8388$ $0.9069$
7: UStyle (Full) $0.9349$ $0.9613$ $0.8260$ $0.8818$ $0.6727$ $0.8389$ $0.8506$ $0.9126$
RMSE 1: UStyle (w/o $\mathcal{L}_{SSIM}$) $10.77$ $9.61$ $10.33$ $9.42$ $10.91$ $10.47$ $10.32$ $9.36$
2: UStyle (w/o $\mathcal{L}_{color}$) $10.65$ $9.45$ $10.19$ $9.32$ $10.47$ $9.90$ $10.25$ $9.30$
3: UStyle (w/o $\mathcal{L}_{fft}$ ) $10.83$ $9.51$ $10.67$ $9.98$ $11.01$ $10.60$ $10.12$ $9.21$
4: UStyle (w/o $\mathcal{L}_{percept}$) $10.23$ $9.30$ $9.81$ $9.25$ $10.76$ $10.23$ $9.98$ $9.09$
5: UStyle (w/o $\mathcal{L}_{clip}$) $10.15$ $9.43$ $9.90$ $9.40$ $10.37$ $10.19$ $9.75$ $9.01$
6: UStyle (w/o $L3$) $10.23$ $9.77$ $9.93$ $9.45$ $10.38$ $10.28$ $9.70$ $9.15$
7: UStyle (Full) $10.08$ $9.39$ $9.89$ $9.35$ $10.31$ $9.50$ $9.70$ $8.96$
@@ V-D Ablation Studies
Additionally, we investigate the contribution of different loss components by evaluating seven configurations of our model: (1-5) UStyle without (w/o) each individual loss component: $\mathcal{L}_{ssim}$, $\mathcal{L}_{color}$, $\mathcal{L}_{fft}$, $\mathcal{L}_{percept}$ and $\mathcal{L}_{clip}$; (6) UStyle without regularization losses, $L3=\mathcal{L}_{ssim}+\mathcal{L}_{color}+\mathcal{L}_{fft}$ (w/o $L3$); and (7) the full UStyle model with all loss components. Table III presents the quantitative comparison across PSNR, SSIM, and RMSE for four underwater style transfer scenarios. The results show that the full UStyle model consistently performs better, demonstrating the effectiveness of depth supervision and loss regularization. From a quantitative perspective, depth supervision leads to a substantial improvement in all three metrics. The inclusion of additional losses further refines the model’s ability to balance structural preservation and perceptual realism. The full UStyle model achieves the highest PSNR and SSIM values while maintaining the lowest RMSE across all scenarios.
To further examine the impact of DA-WCT stylization, we employ Grad-CAM [selvaraju2017grad] visualizations to highlight the regions of interest in our model’s decision-making process. As illustrated in Fig. 8, depth supervision enables UStyle to focus more effectively on salient structures, ensuring spatially coherent style adaptation. The middle row displays activation maps generated with depth guidance, where attention is evenly distributed across foreground and background elements, enhancing structure retention. In contrast, the bottom row corresponding to the model without depth supervision reveals inconsistent attention patterns, leading to fine-detail loss and exaggerated distortions. Notably, DA-WCT ensures that complex structures, such as marine organisms and underwater artifacts, are better preserved by aligning style fusion with content depth information.
@@
TABLE IV: Impact of depth noise on UStyle; we compare depth maps from Monodepth2, UDepth, and DepthAnythingV2 on the UF7D test set, reporting average PSNR ($\uparrow$), SSIM ($\uparrow$), and RMSE ($\downarrow$). The results show that more accurate depth maps yield smoother and more consistent stylization.
(a) Metric Depth Estimator B $\rightarrow$ G G $\rightarrow$ B
PSNR Monodepth2 [godard2019digging] 28.30 28.44
UDepth [yu2022udepth] 28.55 28.66
DepthAnythingV2 [bhat2023depth] 28.73 28.72
SSIM Monodepth2 [godard2019digging] 0.9411 0.8761
UDepth [yu2022udepth] 0.9503 0.8799
DepthAnythingV2 [bhat2023depth] 0.9613 0.8818
RMSE Monodepth2 [godard2019digging] 9.80 9.65
UDepth [yu2022udepth] 9.52 9.40
DepthAnythingV2 [bhat2023depth] 9.39 9.35
(b) Content and style images for the comparison shown on the right.
(c) Stylization with different depth estimators: noisy depth maps from Monodepth2 cause uneven artifacts, UDepth offers moderate improvement, while DepthAnythingV2 generates smoother and coherent stylizations.